OpenStudy (anonymous):

Hi guys.......got one here for ya.... A hardware supplier manufactures three kinds of clamps, types A,B, and C. Production restrictions force it to make 10 more C clamps than the total of the other types and twice as many type B clamps as type A clamps. The shop must produce 310 clamps per day. How many of each type are made per day?

OpenStudy (anonymous):

A+b+c=310. C=a+b+10 b=2a. So. A+2a+A+2a+10=310. 6a+10=310. 6a=300. A=50. And because b=2a. B=100. Because can is a+b+10. C=160

OpenStudy (anonymous):

Hmmmm......

OpenStudy (anonymous):

Oh wow.....you were right!

OpenStudy (anonymous):

You make that seem so easy!